OpenStudy (anonymous):

Simplify square root of negative 100 -10 -10i 10i 10

OpenStudy (dan815):

@abb0t

OpenStudy (abb0t):

omg @dan815 is so lazy.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

pretty sure its C?

OpenStudy (ipwnbunnies):

\[\sqrt{-100} = \sqrt{100} * \sqrt{-1}\]

OpenStudy (anonymous):

so -10?

OpenStudy (ipwnbunnies):

Nope. You know sqrt(100) is 10. Do you know the sqrt of -1? :3

OpenStudy (anonymous):

i?

OpenStudy (ipwnbunnies):

Mhm. c: Now put the two together, or multiply them. OvO

OpenStudy (anonymous):

10i or -10i

OpenStudy (ipwnbunnies):

Which do you think? When written as: \[\sqrt{100} \] They are generally looking for the positive answer of this square root.

OpenStudy (anonymous):

ok 10i so its C

OpenStudy (ipwnbunnies):

Yesh. ^.^

OpenStudy (anonymous):

thank youu!